BJ’s goal is to have $50,000 saved at the end of Year 5. At the end of Year 2, they can add $7,500 to their savings but they want to deposit the remainder they need to reach their goal today, Year 0, as a lump sum deposit. If they can earn 4.5 percent, how much must they deposit today?

rev: 07_04_2017_QC_CS-92838

$33,254.58

$33,108.09

$34,276.34

$34,642.28

$34,912.63
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

50,000=P*(1.045)^5+7500*(1.045)^3

50,000=P*(1.045)^5+8558.745938

P=(50,000-8558.745938)/(1.045)^5

=$33254.58(Approx).